# MegaFusion: Extend Diffusion Models towards Higher-resolution Image Generation without Further Tuning

### v1

@@ B.1 Comparison on crop FID/KID
Following previous work [7], we also calculate the crop FID and crop KID metrics on the generated results of various models to reflect the quality of local patches in the generated images. As depicted in Table 5, previous methods are often limited to specific latent-space models, while our MegaFusion consistently improves the quality of high-resolution image generation in both latent-space and pixel-space models.
@@
B.2 Comparison on CUB-200 Dataset
B.3 More Results of Floyd-MegaFusion
As mentioned above, we also conduct experiments that first generate $128\times 128$ out-of-distribution images, followed by $512\times 512$ high-resolution images on the Floyd model. As depicted in Table 7, under both settings, MegaFusion effectively improves the high-resolution generation capability of Floyd. This demonstrates that MegaFusion can improve the semantic accuracy of high-resolution images at any stage of the generation process.
@@ Appendix C Additional Qualitative Results
C.1 Effects of hyperparameters $\delta$ and $\gamma$
For denoising at the original size, we do not employ dilation. During qualitative experiments for high-resolution generation, we test different $\delta$ values and find that $\delta=2$ is a stable choice for our method and experiment settings, which will not introduce blurriness or semantic deviations. As described in our manuscript, we draw inspiration from simple diffusion [18], which derives the SNR relationship between different resolution images from the mean and variance of pixel distributions. Substituting this into our derived relationship, we obtain that $\gamma=4$. Qualitative experiments also confirm that this is an appropriate choice. Some visualization examples are shown in Figure 5.
@@ C.1 Effects of hyperparameters $\delta$ and $\gamma$
C.2 Text-to-Image Foundation Models
@@
We present more visualization results of higher-resolution image generation using both latent-space and pixel-space text-to-image models in Figure 6 and 7, respectively, to demonstrate the universality and robustness of our proposed method. The visual outcomes explicitly confirm that when pre-trained models fail to scale to higher resolutions, our approach can be universally integrated into existing latent-space and pixel-space diffusion models, improving their capability to synthesize higher-resolution images of megapixels with accurate semantics. Moreover, our further enhanced MegaFusion++ significantly boosts the quality of the generated images, producing sharper and clearer details.
C.3 Comparison to state-of-the-art
@@
To further demonstrate the quality of MegaFusion, we also compare it with existing state-of-the-art high-resolution generation methods. Given that these methods (ScaleCrafter [11] and DemoFusion [7]) are typically limited to specific models, we conduct comparisons on models based on SDXL. The results in Figure 8 indicate that existing methods still face quality degradation and object repetition when generating high-resolution images. In contrast, MegaFusion produces high-quality, semantically accurate high-resolution images. Additionally, as shown in Table 1, MegaFusion is much more efficient than existing approaches.
C.4 Models with additional conditions
@@
We have confirmed that our method is equally applicable to diffusion models with additional input conditions, such as ControlNet [51] with depth maps and IP-Adapter [48] with reference images as extra inputs. As depicted in Figure 9, we further discover that ControlNet [51] with canny edges or human poses as additional conditional inputs also struggle with synthesizing higher-resolution images, and often produce images that are not fidelity to input conditions, with confusing semantics and poor image quality. In contrast, with the assistance of our proposed MegaFusion, our boosted model, ControlNet-MegaFusion is still capable of generating high-quality images of higher resolutions with correct semantics, that are fidelity to conditions. This further illustrates the limitations of existing pre-trained generative models and powerfully demonstrates the effectiveness and universality of our proposed strategy.
C.5 Generation with Arbitrary Aspect Ratios
@@
As previously stated, our MegaFusion can also assist existing pre-trained diffusion models in generating images at arbitrary aspect ratios. As illustrated in Figure 10, 11 and 12, building upon the SDXL [30] model, we offer more qualitative results from SDXL-MegaFusion across various aspect ratios and resolutions, including $1:1$ ($2048\times 2048$), $16:9$ ($1920\times 1080$), $3:4$ ($1536\times 2048$), and $4:3$ ($2048\times 1536$). These impressive visual outcomes further demonstrate the scalability and superiority of our approach.
@@ D.1 Limitations
Since our proposed MegaFusion is a tuning-free approach based on existing latent-space and pixel-space image generation models, it inevitably inherits some of the limitations of current diffusion-based generative models. For example, when with complex textual conditions as inputs, the generated content often struggles to accurately reflect the input prompts, particularly in aspects such as attribute binding and positional control. This may lead to a degradation in synthesis quality during high-resolution generation with MegaFusion. However, better and more powerful backbone models are expected to mitigate this issue, and when combined with MegaFusion, they are likely to produce higher-quality images at higher resolutions with low computational costs.
@@ D.2 Future Work
The striking quantitative results produced by our proposed MegaFusion have confirmed its potential to address the limitations of existing diffusion-based generative models and to improve their capabilities to synthesize high-resolution outcomes. Additionally, we have also observed that existing video generation models encounter significant semantic deviations and quality decline when generating contents that exceed their pre-trained spatial resolution and temporal length. Therefore, we anticipate further applying MegaFusion to current video generation models towards efficient and low-cost higher-resolution longer video content generation. Similarly, MegaFusion also holds the potential for extension to 3D object generation models and models for image and video editing, which are also left for future research.
Figure 6: More qualitative results of applying our MegaFusion to latent-space diffusion model (SDM [35]) for higher-resolution image generation on MS-COCO [24] and commonly used prompts from the Internet.
Figure 7: More qualitative results of applying our MegaFusion to pixel-space diffusion model (Floyd [6]) for higher-resolution image generation on MS-COCO [24] and commonly used prompts from the Internet.
@@
Figure 9: Qualitative results of applying our MegaFusion to ControlNet [51] with canny edges or human poses as extra conditions for higher-resolution image generation with better semantics and fidelity.

### v2

@@ B.1 Comparison on crop FID/KID
Following previous work [9], we also evaluate crop FID and crop KID metrics on the generated results of various models to reflect the quality of local patches in the images. As depicted in Table 5, previous methods are often limited to specific latent-space models, whereas our MegaFusion consistently improves the quality of high-resolution image generation across both latent-space and pixel-space models.
@@
B.2 Comparison on CUB-200 Dataset
B.3 More Results of Floyd-MegaFusion
As mentioned above, we also conduct experiments that first generate $128\times 128$ out-of-distribution images, followed by $512\times 512$ high-resolution images on the Floyd model. As depicted in Table 7, MegaFusion consistently improves the high-resolution generation capability of Floyd under both settings. This demonstrates that MegaFusion can improve the semantic accuracy of high-resolution images at any stage of the generation process.
B.4 Ablation Study of Classifier-free Guidance
As detailed in the implementation details, to ensure a fair comparison and eliminate the impact of classifier-free guidance (CFG) on generation quality and efficiency, we use the default CFG weights from official implementations for all methods and their corresponding MegaFusion-boosted counterparts. To further investigate the impact of CFG on MegaFusion at higher resolutions, we generate 100 images from the MS-COCO dataset using SDM-MegaFusion and SDXL-MegaFusion with varying CFG values, using the same text prompt and random seed as inputs, and evaluate the FID scores against our testset. The results in Figure 5 indicate that classifier-free guidance does affect our high-resolution generation quality, with preliminary findings indicating that $w=7.0$ is a relatively good choice for SDM-MegaFusion and SDXL-MegaFusion.
Figure 5: Ablation study of classifier-free guidance (CFG) weight on SDM-MegaFusion and SDXL-MegaFusion.
@@ Appendix C Additional Qualitative Results
C.1 Evidence Behind the Core idea & intuition
As stated in eDiff-I [1], diffusion models synthesize semantics during early denoising stages and refine image details in later stages. As depicted in Figure 6, we also observe that semantic deviations and object repetitions commonly encountered at higher resolutions primarily stem from incorrect semantics generated during early denoising, leading to irreparable errors. Thus, our intuition and insight here are: perform early denoising at the original resolution to generate accurate semantic information, followed by truncate and relay to continue denoising at higher resolutions, thereby enriching texture details. This enables MegaFusion to produce high-quality, semantically accurate higher-resolution images with lower computational costs, while supporting arbitrary aspect ratios.
Figure 6: Evidence behind our core idea and intuition. For $T=50$ steps of DDIM sampling, we visualize the key stages of the image generation process. For SDM and SDXL, incorrect semantics are generated during the early denoising stages of high-resolution generation, leading to irreparable errors. In contrast, MegaFusion generates accurate semantics and further enriches texture details at higher resolutions. The input text prompts are “A cute black cat” and “A white dog sits on the grass.” For ease of visualization, the images are scaled to the same size.
C.2 Disadvantages of Direct Upsampling
Compared to our MegaFusion for higher-resolution image generation, a more straightforward approach is to directly apply upsampling to images generated by diffusion models. Although simple, this will introduce three potential issues: (i) Direct super-resolution may lead to unrealistic texture details, such as blurring and artifacts, especially at high upsampling factors; (ii) While diffusion-based SR methods can produce more realistic textures via iterative denoising, they often involve significantly higher computational costs and may not support arbitrary aspect ratios; (iii) Most critically, as shown in Figure 7, directly upsampling low-resolution images can stretch and distort content, particularly when generating under non-standard aspect ratios (e.g. $1:4$), diminishing the natural aesthetic of images.
In contrast, MegaFusion seamlessly bridges coarse-to-fine generation processes, efficiently producing accurate semantics at low resolutions and enriching texture details at high resolutions. Leveraging iterative denoising at higher resolutions, it can synthesize aesthetically pleasing high-resolution images even with non-standard aspect ratios.
Figure 7: Analysis of direct upsampling. Using diffusion models to generate images with non-standard aspect ratios directly or via upsampling, may lead to stretching and distortion (e.g., trees on both sides), while MegaFusion effectively mitigates this issue.
C.3 Effects of hyperparameters $\delta$ and $\gamma$
For denoising at the original size, we do not employ dilation. In qualitative experiments for high-resolution generation, we test various $\delta$ values and find that $\delta=2$ is a stable choice under our experimental settings, which will not introduce blurriness or semantic deviations. As described in our manuscript, we draw inspiration from simple diffusion [21], which derives the SNR relationship between images of different resolution based on the mean and variance of pixel distributions. Substituting this into our derived relationship, we obtain that $\gamma=4$. Qualitative experiments also confirm that this is an appropriate choice. Some visualization examples are shown in Figure 8.
@@ C.3 Effects of hyperparameters $\delta$ and $\gamma$
C.4 Ablation Study of Truncation Steps
In the truncate and relay strategy, the number of denoising steps at each stage may also affect generation quality. Our intuition and experience suggest that more denoising steps at lower resolutions improve generation efficiency, while additional steps at higher resolutions enhance texture details. However, conducting a comprehensive evaluation to determine the optimal truncation steps would incur significant computational costs. Therefore, in our implementation, we empirically select truncation steps for each model based on experience, and validate the above conclusions through qualitative experiments, as shown in Figure 9. Considering the trade-off between generation quality and efficiency, we choose denoising steps of $T_{1}=40$, $T_{2}=5$, and $T_{3}=5$ as the default configuration for SDM-MegaFusion.
Figure 9: Qualitative ablation study of truncation steps.
C.5 Text-to-Image Foundation Models
@@
We present more visualizations of higher-resolution image generation using both latent-space and pixel-space text-to-image models in Figure 10 and 11, respectively, to demonstrate the universality and robustness of our proposed method. The visual outcomes explicitly confirm that when pre-trained models fail to scale to higher resolutions, our approach can be universally integrated into existing latent-space and pixel-space diffusion models, improving their capability to synthesize higher-resolution images of megapixels with accurate semantics. Moreover, our further enhanced MegaFusion++ significantly boosts the quality of the generated images, producing sharper and clearer details.
C.6 Compatibility with Transformer-based Models
To further demonstrate the versatility and effectiveness of MegaFusion, we also apply it to the transformer-based (DiT) SD3 [10] model. Since DiT-based methods do not involve convolutions, we boost the model via only the truncate and relay strategy. As presented in Figure 12, SD3 also encounters quality degradation when generating higher-resolution images directly, while our MegaFusion effectively improves its high-resolution generation capabilities.
C.7 Comparison to state-of-the-art
@@
To further evaluate the quality of MegaFusion, we compare it with existing state-of-the-art high-resolution generation methods. Given that these methods (ScaleCrafter [14] and DemoFusion [9]) are typically restricted to specific models, we conduct comparisons on models based on SDXL. The results in Figure 13 indicate that existing methods still face quality degradation and object repetition when generating high-resolution images. In contrast, MegaFusion produces high-quality, semantically accurate high-resolution images, and is much more efficient than existing approaches, as shown in Table 1 of our manuscript.
C.8 Models with additional conditions
@@
We have confirmed that our method is equally applicable to diffusion models with additional input conditions, such as ControlNet [54] with depth maps and IP-Adapter [51] with reference images as extra inputs. As depicted in Figure 14, we further discover that ControlNet with canny edges or human poses as conditional inputs also struggle with synthesizing higher-resolution images, and often produce images that are not fidelity to input conditions, with confusing semantics and poor image quality. In contrast, with the assistance of our proposed MegaFusion, our boosted model, ControlNet-MegaFusion consistently generates high-quality images of higher resolutions with accurate semantics, that are fidelity to conditions.
C.9 Generation with Arbitrary Aspect Ratios
@@
As previously stated, our MegaFusion also enables existing pre-trained diffusion models to generate images at arbitrary aspect ratios. Figure 15, 16 and 17 showcase more qualitative results from SDXL-MegaFusion across various aspect ratios and resolutions, including $1:1$ ($2048\times 2048$), $16:9$ ($1920\times 1080$), $3:4$ ($1536\times 2048$), and $4:3$ ($2048\times 1536$). Moreover, as presented in Figure 18, 19, and 20, we also include visualizations with non-standard aspect ratios, such as $1:4$ ($640\times 2560$), $4:1$ ($2560\times 640$), $1:2$ ($1024\times 2048$), $2:1$ ($2048\times 1024$), $21:9$ ($2016\times 864$), and $9:21$ ($864\times 2016$). These impressive outcomes further demonstrate the scalability and superiority of our approach.
C.10 Compatibility with LoRA
To further illustrate the versatility and broad applicability of MegaFusion, we apply it to SDM and SDXL models using LoRA from the open-source community for personalized higher-resolution image generation. As depicted in Figure 21, MegaFusion can seamlessly integrate with various LoRAs of SDM and SDXL, demonstrating significant potential for artistic and commercial applications.
@@ D.1 Limitations
Since our proposed MegaFusion is a tuning-free approach built on existing latent-space and pixel-space image generation models, it inevitably inherits some limitations of current diffusion-based generative models. For example, when handling complex textual conditions, the generated content often struggles to accurately reflect input prompts, particularly in aspects such as attribute binding and positional control. This may lead to degraded synthesis quality during high-resolution generation with MegaFusion. However, more powerful backbone models are expected to mitigate this issue, and when combined with MegaFusion, they are likely to produce higher-quality images at higher resolutions with low computational costs.
@@ D.2 Future Work
The striking quantitative results produced by MegaFusion have confirmed its potential to overcome the limitations of existing diffusion-based generative models and to improve their capabilities to synthesize high-resolution outcomes. Additionally, we have observed that existing video generation models encounter significant semantic deviations and quality degradation when generating content beyond their pre-trained spatial resolution and temporal length. Therefore, we anticipate further applying MegaFusion to current video generation models towards efficient, low-cost, higher-resolution, and longer video content generation. Similarly, MegaFusion also holds the potential for extension to 3D generation models and models for image and video editing, which are also left for future exploration.
Figure 10: More qualitative results of applying our MegaFusion to latent-space diffusion model (SDM [37]) for higher-resolution ($1024\times 1024$) image generation on MS-COCO [26] and commonly used prompts from the Internet.
Figure 11: More qualitative results of applying our MegaFusion to pixel-space diffusion model (Floyd [8]) for higher-resolution ($512\times 512$) image generation on MS-COCO [26] and commonly used prompts from the Internet.
Figure 12: Qualitative results of applying our MegaFusion to latent-space diffusion model (SD3 [10]) for higher-resolution ($2048\times 2048$) image generation on MS-COCO [26] and commonly used prompts from the Internet.
@@
Figure 14: Qualitative results of applying our MegaFusion to ControlNet [54] with canny edges or human poses as extra conditions for higher-resolution ($1024\times 1024$) image generation with better semantics and fidelity.
@@
Figure 18: textbfMore qualitative results of applying our MegaFusion to SDXL [32] model for higher-resolution image generation with various non-standard aspect ratios and resolutions.
Figure 19: More qualitative results of applying our MegaFusion to SDXL [32] model for higher-resolution image generation with various non-standard aspect ratios and resolutions.
Figure 20: More qualitative results of applying our MegaFusion to SDXL [32] model for higher-resolution image generation with various non-standard aspect ratios and resolutions.
Figure 21: Qualitative results of applying MegaFusion to high-resolution image generation with LoRA-integrated SDM and SDXL. Similarly, SDM and SDXL integrated with LoRA also face common challenges like semantic deviations and object repetitions in high-resolution generation, while MegaFusion effectively addresses these challenges.